Find the variation constant and an equation of variation where y varies directly as x and y80 when x=10
THe variation constant is K=
The equation of variation is y=

Answer by bucky(2189)

If y varies directly as x, then the variation equation is of the form:
.
y = K*x
.
You are told that y = 80 when x = 10. Substitute these two values into the equation and
you get:
.
80 = K*10
.
Solve for K by dividing both sides of the equation by 10 to get:
.
80/10 = K
.
The left side reduces to 8 so you have:
.
8 = K
.
Now you can return to the form of the "direct variation" equation and substitute 8 for k to
get the equation:
.
y = 8*x
.
And there you are ...
.
the variation constant is K = 8
.
and the equation of variation is y = 8*x
.
Hope this helps you to understand the problem. The key to these problems is knowing that
the equation of direct variation is of the form:
.
y = K*x
.
And the equation of inverse variation is of the form:
.
y = K/x
.